# LS-category and topological complexity of real torus manifolds and Dold manifolds of real torus type

### 12 Jan 2024 v1

@@ Abstract.
The real torus manifolds are a generalization of small covers and generalized real Bott manifolds. We compute the LS-category of these manifolds under some constraints and obtain sharp bounds on their topological complexities. We obtain a simplified description of their cohomology ring and discuss a relation on the cup-product of its generators. We obtain the sharp bounds on their zero-divisors-cup-lengths. We improve the dimensional upper bound on their topological complexity. We also show that under certain hypotheses, the topological complexity of real torus manifolds of dimension $n$ is either $2n$ or $12n+$ 1. We compute the $\mathbb{Z}_{2}$-equivariant LS-category of any small cover when the $\mathbb{Z}_{2}$-fixed points are path connected. We then compute the LS-category of Dold manifolds of real torus type and obtain sharp bounds on their topological complexity. In the end, we obtain sharp bounds on the symmetric topological complexity of a class of these manifolds.
@@ 1. Introduction
Farber [16] introduced the concept of topological complexity to study the robot motion planning problem through the topological lens. For a topological space $X$, the topological complexity $\mathrm{TC}(X)$ is a numerical homotopy invariant. Let $X$ be a path-connected space and $PX$ be the space of all paths in $X$ equipped with a compact open topology. Let $\gamma\colon[0,1]\to X$ be any path in $X$. Then there is a fibration $\pi\colon PX\to X\times X$ defined by $\pi(\gamma)=(\gamma(0),\gamma(1))$. The topological complexity of $X$ is the smallest $k$ for which $X\times X$ admit an open cover $U_{1},\dots,U_{k}$, such that there exist continuous sections of $\pi$ on $U_{i}$ for $1\leq i\leq k$. In general, determining the exact value of $\mathrm{TC}(X)$ is a hard problem even for the spaces having nice CW-complex structures like Dold manifolds. Since Farber first introduced this invariant in $2003$, numerous methods have been created to determine bounds on $\mathrm{TC}(X)$, and the precise value of this invariant has been calculated for several spaces, see for example [3], [13], [14], [16], [18], [20], [22], [23]. This invariant is closely related to the well-known invariant, the Lusternik-Schnirelmann category (in short, LS-category) [26]. For a path connected space $X$, the LS-category of $X$ is denoted $\mathrm{cat}(X)$. The $\mathrm{cat}(X)$ is the smallest integer $r$ such that $X$ can be covered by $r$ open subsets $V_{1},\dots,V_{r}$ with inclusion $V_{i}\hookrightarrow{}X$ is null-homotopic for $1\leq i\leq r$. In particular, it was shown in [7] that,
@@
The cohomological methods have been used to obtain some lower bounds on $\mathrm{TC}(X)$ and $\mathrm{cat}(X)$. We describe them here briefly. Let $R$ be a commutative ring with unity and $X$ be a path-connected topological space with its cohomology ring $H^{\ast}(X;R)$. The cup-length of $X$ over $R$ is the maximal integer $\ell$ such that there exists $x_{i}\in H^{\ast}(X;R)$ for $i=1,\ldots,\ell$ satisfying $\prod_{i=1}^{\ell}x_{i}\neq 0$. We denote this number by $\mathrm{cl}_{R}(X)$. It was proved in [7, Proposition 1.5] that the number gives a lower bound for the $\mathrm{cat}(X)$. Let
@@
be the map induced by the cup product. Then the zero-divisors-cup-length of $X$ with respect to the coefficient ring $R$ is defined as the maximal integer $k$ such that there exist cohomology classes ${u_{i}\in H^{\ast}(X;R)\otimes H^{\ast}(X;R)}$ satisfying $\cup(u_{i})=0$ for all $1\leq i\leq k$ and $\prod_{i=1}^{k}u_{i}\neq 0$. This integer is denoted by $\mathrm{zcl}_{R}(X)$. It was proved in [16, Theorem 7] that $\mathrm{zcl}_{R}(X)$ gives a lower bound for the $\mathrm{TC}(X)$. That is, we have the following;
@@
A well-studied manifolds are generalized Bott manifolds. However, little is known about their real counterpart. Briefly, A generalized real Bott tower of height $m$ is a sequence
@@
of manifolds $B_{j}=\mathbb{P}(\underline{\mathbb{R}}\oplus E_{j}^{(1)}\oplus\cdots\oplus E_{j}^{(n_{j})})$, where $\underline{\mathbb{R}}$ is the trivial line bundle over $B_{j-1}$, $E_{j}^{(i)}$ is a real line bundle over $B_{j-1}$ for $i=1,\dots,n_{j}$, and $j=1,\dots,m$. Here $\mathbb{P}(\cdot)$ denotes the projectivization. The space $B_{j}$ is called a $j$-th stage generalized real Bott manifold. Note that the fibre of the map $\pi_{j}$ is $\mathbb{R}P^{n_{j}}$ for $j=1,\ldots,m$. In particular, $B_{1}=\mathbb{R}P^{n_{1}}$. Also, when $n_{j}=1$ for every $j\in\{1,\ldots,m\}$, then $B_{j}$ is called an $m$th-stage real Bott manifold. Interestingly, generalized real Bott manifolds are small covers. We note that the small covers are central objects in Toric Topology introduced by Davis and Januszkiewicz [11]. They are smooth manifolds with effective locally standard real torus actions and have rich combinatorial properties. The concept of small covers has been extended in [25] where the author called them $2$-torus manifolds. If the real torus action on a $2$-torus manifold is locally standard and the orbit space is a nice manifold with corners, we may call the $2$-torus manifold a ‘real torus manifold’. Following [25], one can get a combinatorial pair $(P,\lambda)$ from a real torus manifold where $P$ is the orbit space which is a nice manifold with corners and $\lambda$ is called a $\mathbb{Z}_{2}$-characteristic function on $P$.
@@
The other class of manifolds we are interested in are called generalized projective product spaces. These manifolds were recently introduced by S. Sarkar and P. Zvengrowski in [29], extending the concept of (generalized) Dold manifolds [12, 27] and projective product spaces [9]. Let $N$ be a manifold equipped with a free $\mathbb{Z}_{2}$-action and $M$ a $\mathbb{Z}_{2}$-manifold. Then the diagonal $\mathbb{Z}_{2}$-action on the product $M\times N$ is free. Then the orbit space $D(M,N):=(M\times N)/\mathbb{Z}_{2}$ is a manifold. In particular, if $M$ is a real torus manifold, we call $D(M,N)$ a Dold manifold of real torus type. Several properties of these manifolds remain to explore for many $M$ and $N$. The goal of this article is to compute the LS-category and topological complexity of real torus manifolds (which contain generalized real Bott manifolds and small covers) and Dold manifolds of real torus type. We explain the details of this in the organization.
@@
In Section 2, we recall a few basic topological properties of real torus manifolds. we prove that the LS-category of a real torus manifold $M$ is $\dim(M)+1$ if the orbit space contains a boundary of a simple polytope, see Theorem 2.5. If $M$ is an $n$-dimensional real torus manifold such that a facet of the orbit space is an $(n-1)$-simplex, then the lower bound on the topological complexity of $M$ can be given by the zero-divisors-cup-length of $\mathbb{R}P^{n-1}$, see Theorem 2.7. In addition, if the $\mathbb{Z}_{2}$-characteristic function $\lambda$ associated with $M$ satisfies certain conditions, then the lower bound on the topological complexity of $M$ can be given by the zero-divisors-cup-length of $\mathbb{R}P^{n}$, see Theorem 2.7. Then, we show that a certain product of a generating set of $H^{*}(M(P,\lambda);\mathbb{Z}_{2})$ is non-zero if $P$ is a finite product of simplices, see Lemma 2.12. As a result, we use cohomological methods to obtain a sharp lower bound on the topological complexity of generalized real Bott manifolds, see Theorem 2.13 and Corollary 2.14. Then we prove that if the product of simplices contains $k$-many odd dimensional simplices, then the dimensional upper bound on the topological complexity of generalized real Bott manifolds can be improved by $k$, see Theorem 2.15.
@@
In Section 3, we compute the $\mathbb{Z}_{2}$-equivariant LS-category of small covers when the fixed point sets are path connected, see Theorem 3.2. For the other kinds of $\mathbb{Z}_{2}$-actions, when fixed point sets are not connected, we show that the $\mathbb{Z}_{2}$-equivariant category of small covers could be much greater than the dimension plus one, see Theorem 3.3. In general, it remains lesser than the number of vertices of the corresponding polytope. Example 3.4 observes that, in certain cases, it can be exactly the number of vertices of the polytope. Then we compute the LS-category of Dold manifolds of real torus type and obtain sharp bounds on their topological complexity, see Theorem 3.7 and Theorem 3.8. In many specific cases, the topological complexity of these manifolds is either $a-1$, $a$, or $a+1$ for some $a\in\mathbb{N}$, see Corollary 3.9.
In Section 4, first we recall the definition and a few properties of symmetric topological complexities. Then, we obtain sharp bounds on the symmetric topological complexity of a class of real torus manifolds and Dold manifolds of real torus type (see Theorem 4.4, Theorem 4.5 and Theorem 4.6). Moreover, in Theorem 4.5, we compute the exact value of symmetric topological complexity of many real torus manifolds.
@@ 2. LS-category and topological complexity of real torus manifolds
In this section, we recall the concepts of real torus manifolds and generalized real Bott manifolds. We give a nice presentation of the cohomology ring of generalized real Bott manifolds with $\mathbb{Z}_{2}$-coefficients. Then, we compute some lower and upper bounds for the topological complexity of these manifolds. We begin by stating the concept of real torus manifolds.
We recall the definition of a nice manifold with corners following [10] and [2]. The codimension function $c$ associates a point $x=(x_{1},\dots,x_{n})\in(\mathbb{R}_{\geq 0})^{n}$ to the number of $x_{i}$ which are zero. An $n$-dimensional manifold with corners is a Hausdorff, second-countable topological space together with a maximal atlas of local charts onto open subsets of $(\mathbb{R}_{\geq 0})^{n}$ such that the overlapping maps are homeomorphisms that preserve codimension function. That means the codimension function is a well-defined map from a manifold with corners $P$ to non-negative integers. A connected component of $c^{-1}(m)$ is called an open face of $P$. The closure of an open face is called a face. We can define the dimension of a face by subtracting its codimension from $n$. In an $n$-dimensional nice manifold with corners, a face of dimension zero and $n-1$ are called a vertex and a facet. The vertex set and the facet set of $P$ are denoted by $V(P)$ and $\mathcal{F}(P)$, respectively.
A manifold with corners $P$ is called nice if, for every $p\in P$ with $c(p)=2$, the number of facets of $P$ which contains $p$ is also $2$. Therefore, a codimension-$k$ face of a nice manifold with corners $P$ is a connected component of the intersection of the unique collection of $k$ many facets of $P$.
An easy example of a manifold with corners is a convex polytope (a convex hull of finitely many points in $\mathbb{R}^{n}$ for some $n\in\mathbb{Z}_{\geq 0}$). If a convex polytope is nice as in the above sense, then it is called a simple polytope i.e., at each vertex exactly $n$ many facets intersect.
Consider $\mathbb{R}^{n}$ with standard $\mathbb{Z}_{2}$-action induced by the reflections on coordinate hyperplanes. A smooth action of $\mathbb{Z}_{2}^{n}$ on an $n$-dimensional smooth manifold $M$ is called locally standard, if $x\in M$ then it has a $\mathbb{Z}_{2}^{n}$-invariant open neighbourhood $U_{x}$ and a diffeomorphism $\phi\colon U_{x}\to V$, where $V$ is $\mathbb{Z}_{2}^{n}$-invariant open set in $\mathbb{R}^{n}$, and an isomorphism $\delta_{x}:\mathbb{Z}_{2}^{n}\to\mathbb{Z}_{2}^{n}$ such that $\phi(t\cdot x)=\delta_{x}(t)\cdot\phi(x)$ for all $t\in\mathbb{Z}_{2}^{n}$ and $x\in U_{x}$.
@@
The identification space
$M^{n}(P,\lambda):=(P\times\mathbb{Z}_{2}^{n})/\sim$
has an $n$-dimensional manifold structure with a natural locally standard $\mathbb{Z}_{2}^{n}$-action induced by the group operation on the second factor of $P\times\mathbb{Z}_{2}^{n}$. Thus $M^{n}(P,\lambda)$ is a real torus manifold over $P$ (see [11] for details). The projection onto the first factor gives the orbit map
@@
where $[p,g]_{\sim}$ is the equivalence class of $(p,g)$.
@@
[…]
Define $\lambda\colon\mathcal{F}(Q)\to\mathbb{Z}_{2}^{n}$ by $\lambda(F_{i})=(1,\dots,1,-1,1,\dots,1)$, where $-1$ is at the $i$-th component. Then $M^{n}(Q,\lambda)$ is a real torus manifold homeomorphic to $S^{n}$, which is not a small cover.
Let $M$ be a real torus manifold over a nice manifold with corners $P$ of dimension $n$, and $\mathfrak{q}\colon M\to P$ be the orbit map. Let $\mathcal{F}(P)=\{F_{1},\ldots,F_{r}\}$. Then the isotropy subgroup of $\mathfrak{q}^{-1}(F_{i})$ is a subgroup of $\mathbb{Z}_{2}^{n}$ and isomorphic to $\mathbb{Z}_{2}$. This isotropy subgroup is uniquely determined by $\lambda_{i}\in\mathbb{Z}_{2}^{n}\setminus\{(1,\ldots,1)\}$ for $i=1,\ldots,m$. Thus, we have an assignment $\lambda\colon\mathcal{F}(P)\to\mathbb{Z}_{2}^{n}$ satisfying Definition 2.2. Therefore, by [25, Lemma 3.1], we have the following.
@@
Let $M$ be an $n$-dimensional real torus manifold over $P$. If a connected component of $\partial P$ is a boundary of an $n$-dimensional simple polytope $Q$, then $\mathrm{cat}(M)=n+1$.
@@
Let $v$ be a vertex in $\partial Q\subseteq\partial P$. So there exist unique facets $F_{1},\ldots,F_{n}$ of $P$ such that $v=F_{1}\cap\cdots\cap F_{n}$ and $F_{i_{1}}\cap\cdots\cap F_{i_{k}}$ is an $(n-k)$-dimensional simple polytope for any $\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$. Let $\mathfrak{q}\colon M\to P$ be the orbit map and $x_{i}$ the Poincare dual of $\mathfrak{q}^{-1}(F_{i})$. It follows from the given hypotheses on $P$, that the intersection $F_{1}\cap\cdots\cap F_{n}$ is transversal. Thus, the product $x_{1}\cdots x_{n}$ of cohomology classes is nonzero in $H^{*}(M;\mathbb{Z}_{2})$. Therefore, the result follows from the cup-length lower bound (1) and the dimensional upper bound as given in [7, Theorem 1.7]. ∎
@@
[…]
@@
Precisely, if $F_{i}$ be the $i$-th facet of $P$ and $x_{i}\in H^{1}(M;\mathbb{Z}_{2})$ corresponds to $F_{i}$, then $\iota^{*}(x_{i})\in H^{1}(\mathbb{R}P^{n-1};\mathbb{Z}_{2})$ corresponds to $F_{i}\cap\Delta^{n-1}$. Note that $F_{i}\cap\Delta^{n-1}$ is empty then $\iota^{*}(x_{i})=0$. Since $\iota^{*}$ is surjective, we have
@@
Thus, the left inequality of (4) follows. The right inequality of (4) follows from [16, Theorem 4].
@@
Let $M$ be an $n$-dimensional real torus manifold over $P$ with $n>1$ and $\Delta^{n-1}$ a facet of $P$. Suppose $F_{1},\ldots,F_{n}$ be the facets of $P$ such that $\Delta^{n-1}\cap F_{i}\neq\emptyset$ and $\det[\lambda(F_{1})\cdots\lambda(F_{n})]=1$. Then $\mathrm{cat}(M)=n+1$, and
@@
In particular, if $n=2^{s}$, then $12^{s+1}\leq\mathrm{TC}(M)\leq 2^{s+1}+$ 1 .
@@
We can have $M=M^{\prime}\#\mathbb{R}P^{n}$ following the discussion in Subsection 1.11 in [11]. Then $H^{*}(\mathbb{R}P^{n};\mathbb{Z}_{2})$ is a subring of $H^{*}(M;\mathbb{Z}_{2})$. Then $n+1=\mathrm{cl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})\leq\mathrm{cl}_{\mathbb{Z}_{2}}(M)\leq n+1$, and the left inequality of (5) follows from the observation $\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})\leq\mathrm{zcl}_{\mathbb{Z}_{2}}(M^{n}(P,\lambda))$. The right inequality of (5) follows from [16, Theorem 4].

### 04 Apr 2026 v2

@@ Abstract.
The real torus manifolds are a generalization of small covers, and the Dold manifolds of real torus type are a class of non-trivial fibre bundles over the projective product spaces with real torus manifolds as fibres. In this paper, first, we compute the LS-category of these two types of manifolds and obtain sharp bounds on their topological complexities. We show that under certain hypotheses, the topological complexities of real torus manifolds of dimension $n$ are either $2n$ or $2n+1$. We figure out tight bounds for the topological complexity of generalized real Bott manifolds, and in many cases, the difference between these upper and lower bounds is less than 5. We compute the $\mathbb{Z}_{2}$-equivariant LS-category of small covers when the $\mathbb{Z}_{2}$-fixed points are path connected. In the end, we study the symmetric topological complexity of the above-mentioned manifolds and obtain exact values for infinitely many cases.
@@ 1. Introduction
Farber [17] introduced the concept of topological complexity to study the robot motion planning problem through the topological lens. For a topological space $X$, the topological complexity $\mathrm{TC}(X)$ is a numerical homotopy invariant. Let $X$ be a path-connected space and $PX$ be the space of all paths in $X$ equipped with the compact open topology. Let $\gamma\colon[0,1]\to X$ be any path in $X$. Then there is a fibration $\pi\colon PX\to X\times X$ defined by $\pi(\gamma)=(\gamma(0),\gamma(1))$. The topological complexity of $X$ is the smallest $k$ for which $X\times X$ admits an open cover $U_{1},\dots,U_{k}$, such that there exist continuous sections of $\pi$ on $U_{i}$ for $1\leq i\leq k$. In general, determining the exact value of $\mathrm{TC}(X)$ is a hard problem, even for spaces with nice CW-complex structures like Dold manifolds. Since Farber first introduced this invariant in $2003$, numerous methods have been created to determine bounds on $\mathrm{TC}(X)$, and the precise value of this invariant has been calculated for several spaces, see for example [12], [13], [17], [16], [19], [20], [21]. This invariant is closely related to the well-known invariant, the Lusternik-Schnirelmann category (in short, LS-category) [25]. For a path connected space $X$, the LS-category of $X$ is denoted $\mathrm{cat}(X)$. The $\mathrm{cat}(X)$ is the smallest integer $r$ such that $X$ can be covered by $r$ open subsets $V_{1},\dots,V_{r}$ for which the inclusion $V_{i}\hookrightarrow{}X$ is null-homotopic for $1\leq i\leq r$. In particular, it was shown in [6] that,
@@
The cohomological methods have been used to obtain some lower bounds on $\mathrm{TC}(X)$ and $\mathrm{cat}(X)$. We describe them here briefly. Let $R$ be a commutative ring with unity and $X$ be a path-connected topological space with its cohomology ring $H^{\ast}(X;R)$. The cup-length of $X$ over $R$ is the maximal integer $\ell$ such that there exists $x_{i}\in{\tilde{H}}^{\ast}(X;R)$ for $i=1,\ldots,\ell$ satisfying $\prod_{i=1}^{\ell}x_{i}\neq 0$. We denote this number by $\mathrm{cl}_{R}(X)$. The number gives a lower bound for the $\mathrm{cat}(X)$, see [6, Proposition 1.5]. Let
@@
be the map induced by the cup product. Then the zero-divisors-cup-length of $X$ with respect to the coefficient ring $R$ is defined as the maximal integer $k$ such that there exist cohomology classes $u_{i}\in H^{\ast}(X;R)\otimes H^{\ast}(X;R)$} satisfying $\cup(u_{i})=0$ for all $1\leq i\leq k$ and $\prod_{i=1}^{k}u_{i}\neq 0$. This integer is denoted by $\mathrm{zcl}_{R}(X)$. It was proved in [17, Theorem 7] that $\mathrm{zcl}_{R}(X)$ gives a lower bound for the $\mathrm{TC}(X)$. That is, we have the following;
@@
Several properties of generalized Bott manifolds have been studied. However, little is known about their real counterpart. Briefly, A generalized real Bott tower of height $m$ is a sequence
@@
of manifolds $B_{j}=\mathbb{P}(\underline{\mathbb{R}}\oplus E_{j}^{(1)}\oplus\cdots\oplus E_{j}^{(n_{j})})$, where $\underline{\mathbb{R}}$ is the trivial line bundle over $B_{j-1}$, $E_{j}^{(i)}$ is a real line bundle over $B_{j-1}$ for $i=1,\dots,n_{j}$, and $j=1,\dots,m$. Here $\mathbb{P}(\cdot)$ denotes the projectivization. The space $B_{j}$ is called a $j$-th stage generalized real Bott manifold. Note that the fiber of the map $\pi_{j}$ is $\mathbb{R}P^{n_{j}}$ for $j=1,\ldots,m$. In particular, $B_{1}=\mathbb{R}P^{n_{1}}$. Also, when $n_{j}=1$ for every $j\in\{1,\ldots,m\}$, then $B_{j}$ is called an $j$th-stage real Bott manifold. Interestingly, generalized real Bott manifolds are small covers. We note that the small covers are central objects in Toric Topology introduced by Davis and Januszkiewicz [10]. They are smooth manifolds with effective locally standard real torus actions and have rich combinatorial properties. The concept of small covers has been extended in [24] where the author called them $2$-torus manifolds. If the real torus action on a $2$-torus manifold is locally standard and the orbit space is a nice manifold with corners, we may call the $2$-torus manifold a ‘real torus manifold’. Following [24], one can get a combinatorial pair $(P,\lambda)$ from a real torus manifold where $P$ is the orbit space which is a nice manifold with corners, and $\lambda$ is called a $\mathbb{Z}_{2}$-characteristic function on $P$. We refer to [10, 24, 23] for several topological information on small covers and real torus manifolds. We note that the article [3] calculated the LS-category of a small cover and studied the topological complexity of a specific subclass of generalized real Bott manifolds.
Farber, Tabachnikov and Yuzvinsky [16] showed a remarkable relation between the topological complexity of a real projective space and its immersion dimension. The topological complexity of any generalized real Bott manifolds, which are smooth projective real toric manifolds, is unknown until now. We note that a real torus manifold is a finite quotient of a real moment angle manifold that has several similar topological properties with the same dimensional sphere. Therefore, we think that the following question is interesting and challenging.
Question 1.
Is the topological complexity of a real torus manifold $M$ equal to its immersion dimension?
@@
The other class of manifolds we are interested in are called generalized projective product spaces. These manifolds were recently introduced by Sarkar and Zvengrowski in [27], extending the concept of (generalized) Dold manifolds [11, 26] and projective product spaces [8]. Let $N$ be a manifold equipped with a free $\mathbb{Z}_{2}$-action and $M$ a $\mathbb{Z}_{2}$-manifold. Then the diagonal $\mathbb{Z}_{2}$-action on the product $M\times N$ is free. Then the orbit space $D(M,N):=(M\times N)/\mathbb{Z}_{2}$ is a manifold. In particular, if $M$ is a real torus manifold, we call $D(M,N)$ a Dold manifold of real torus type. Several properties of these manifolds remain to explore for many $M$ and $N$. The goal of this article is to compute the LS-category and topological complexity of real torus manifolds (which contain generalized real Bott manifolds and small covers) and Dold manifolds of real torus type. We explain the details of this in the organization.
@@
In Section 2, we recall some basic topological properties of real torus manifolds. We prove that the LS-category of a real torus manifold $M$ is $\dim(M)+1$ if the orbit space contains a boundary of a simple polytope, see Theorem˜2.5. If $M$ is an $n$-dimensional real torus manifold such that a facet of the orbit space is an (n1)(n-1)-simplex, then the lower bound on the topological complexity of $M$ can be given by the zero-divisors-cup-length of $\mathbb{R}P^{n-1}$, see Theorem˜2.7. In addition, if the $\mathbb{Z}_{2}$-characteristic function $\lambda$ associated with $M$ satisfies certain conditions, then the lower bound on the topological complexity of $M$ can be given by the zero-divisors-cup-length of $\mathbb{R}P^{n}$, see Theorem˜2.7. Then, we show that a certain product of a generating set of $H^{*}(M(P,\lambda);\mathbb{Z}_{2})$ is non-zero if $P$ is a finite product of simplices, see Lemma˜2.11. As a result, we use cohomological methods to obtain a sharp lower bound on the topological complexity of generalized real Bott manifolds, see Theorem˜2.12 and Corollary˜2.13. Then we prove that if the product of simplices contains $k$-many odd dimensional simplices, then the dimensional upper bound on the topological complexity of generalized real Bott manifolds can be improved by $k$, see Theorem˜2.14.
@@
In Section˜3, we compute the $\mathbb{Z}_{2}$-equivariant LS-category of small covers when the fixed point sets are path connected, see Theorem˜3.2. For the other kinds of $\mathbb{Z}_{2}$-actions, when fixed point sets are not connected, we show that the $\mathbb{Z}_{2}$-equivariant category of small covers could be much greater than the dimension plus one, see Theorem˜3.3. In general, it remains less than the number of vertices of the corresponding polytope. Example˜3.4 observes that, in certain cases, it can be exactly the number of vertices of the polytope. Then we compute the LS-category of Dold manifolds of real torus type and obtain sharp bounds on their topological complexity, see Theorem˜3.7 and Theorem˜3.8. In many specific cases, the topological complexity of these manifolds is either a1a-1, $a$, or $a+1$ for some $a\in\mathbb{N}$, see Corollary˜3.9. As an application, we obtain sharp bounds on the symmetric topological complexity of a class of real torus manifolds and Dold manifolds of real torus type (see Theorem˜3.14, Theorem˜3.15, and Theorem˜3.16). Moreover, in Theorem˜3.15, we compute the exact value of the symmetric topological complexity of many real torus manifolds.
@@ 2. LS-category and topological complexity of real torus manifolds
In this section, we recall the concepts of real torus manifolds and generalized real Bott manifolds. We compute the LS-category of a class of real torus manifolds containing small covers and give some tight bounds for the topological complexity of certain real torus manifolds. We describe a nice presentation of the cohomology ring of generalized real Bott manifolds with $\mathbb{Z}_{2}$-coefficients. Then, we obtain sharp bounds for the topological complexity of these manifolds.
We recall the definition of a nice manifold with corners following [9] and [2]. The codimension function $c$ associates to a point $x=(x_{1},\dots,x_{n})\in(\mathbb{R}_{\geq 0})^{n}$ the number of $x_{i}$ which are zero. An $n$-dimensional manifold with corners is a Hausdorff, second-countable topological space together with a maximal atlas of local charts onto open subsets of $(\mathbb{R}_{\geq 0})^{n}$ such that the overlapping maps are homeomorphisms that preserve codimension function. In an $n$-dimensional manifold with corners, a face of codimension $n$ and $1$ are called a vertex and a facet, respectively. The vertex set and the facet set of $P$ are denoted by $V(P)$ and $\mathcal{F}(P)$, respectively.
A manifold with corners $P$ is called nice if, for every $p\in P$ with $c(p)=2$, the number of facets of $P$ that contain $p$ is also $2$. Therefore, a codimension-$k$ face of a nice manifold with corners $P$ is a connected component of the intersection of the unique collection of $k$ many facets of $P$. In particular, if a convex polytope is nice, as in the above sense, it is called a simple polytope.
Consider $\mathbb{R}^{n}$ with standard $\mathbb{Z}_{2}$-action induced by the reflections on coordinate hyperplanes. A smooth action of $\mathbb{Z}_{2}^{n}$ on an $n$-dimensional smooth manifold $M$ is called locally standard, if for all $x\in M$ there exists a $\mathbb{Z}_{2}^{n}$-invariant open neighborhood $U_{x}$ and a diffeomorphism $\phi\colon U_{x}\to V$, where $V$ is $\mathbb{Z}_{2}^{n}$-invariant open set in $\mathbb{R}^{n}$, and an isomorphism $\delta_{x}:\mathbb{Z}_{2}^{n}\to\mathbb{Z}_{2}^{n}$ such that $\phi(t\cdot x)=\delta_{x}(t)\cdot\phi(x)$ for all $t\in\mathbb{Z}_{2}^{n}$ and $x\in U_{x}$.
@@
The identification space $M^{n}(P,\lambda):=(P\times\mathbb{Z}_{2}^{n})/\sim$ has an $n$-dimensional manifold structure with a natural locally standard $\mathbb{Z}_{2}^{n}$-action induced by the group operation on the second factor of $P\times\mathbb{Z}_{2}^{n}$. Thus $M^{n}(P,\lambda)$ is a real torus manifold over $P$ (see [10] for details). The projection onto the first factor gives the orbit map
@@
where $p,g]_{\sim}$ is the equivalence class of $(p,g)$.
@@
Let $Q=\{(x_{1},\ldots,x_{n},x)\in\mathbb{R}^{n+1}~|~\sum_{1}^{n}x_{i}^{2}+x^{2}=1,x_{i}\geq 0~\mbox{for}~i=1,\ldots,n\}$ for $n\geq 2$. Then $Q$ is a nice manifold with corners homeomorphic to a suspension on (n1)(n-1)-simplex. Note that $Q$ is not a simple polytope. The facet set of $Q$ is $\mathcal{F}(Q)=\{F_{1},\dots,F_{n}\}$, where
$F_{i}:=\{(x_{1},\ldots,x_{n},x)\in Q~|~x_{i}=0~\mbox{for}~i=1,\ldots,n\}.$
Define $\lambda\colon\mathcal{F}(Q)\to\mathbb{Z}_{2}^{n}$ by $\lambda(F_{i})=(1,\dots,1,-1,1,\dots,1)$, where 1-1 is at the $i$-th component. Then $M^{n}(Q,\lambda)$ is a real torus manifold homeomorphic to $S^{n}$, which is not a small cover.
Let $M$ be a real torus manifold over a nice manifold with corners $P$ of dimension $n$, and $\mathfrak{q}\colon M\to P$ be the orbit map. Let $\mathcal{F}(P)=\{F_{1},\ldots,F_{r}\}$. Then the isotropy subgroup of $\mathfrak{q}^{-1}(F_{i})$ is a subgroup of $\mathbb{Z}_{2}^{n}$ and isomorphic to $\mathbb{Z}_{2}$. This isotropy subgroup is uniquely determined by $\lambda_{i}\in\mathbb{Z}_{2}^{n}\setminus\{(1,\ldots,1)\}$ for $i=1,\ldots,m$. Thus, we have an assignment $\lambda\colon\mathcal{F}(P)\to\mathbb{Z}_{2}^{n}$ satisfying Definition 2.2. Therefore, by [24, Lemma 3.1], we have the following.
@@
Let $M$ be an $n$-dimensional real torus manifold over $P$. If a connected component of $\partial P$ is a boundary of an $n$-dimensional simple polytope $Q$, then
$\mathrm{cat}(M)=n+1.$
@@
Let $v$ be a vertex in $\partial Q\subseteq\partial P$. So there exist unique facets $F_{1},\ldots,F_{n}$ of $P$ such that $v=F_{1}\cap\cdots\cap F_{n}$ and $F_{i_{1}}\cap\cdots\cap F_{i_{k}}$ is an (nk)(n-k)-dimensional simple polytope for any $\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$. Let $\mathfrak{q}\colon M\to P$ be the orbit map and $x_{i}$ the Poincare dual of $\mathfrak{q}^{-1}(F_{i})$. It follows from the given hypotheses on $P$, that the intersection $F_{1}\cap\cdots\cap F_{n}$ is transversal. Thus, the product $x_{1}\cdots x_{n}$ of cohomology classes is nonzero in $H^{*}(M;\mathbb{Z}_{2})$. Therefore, the result follows from the cup-length lower bound (1) and the dimensional upper bound as given in [6, Theorem 1.7]. ∎
@@
$n\leq\mathrm{cat}(M)\leq n+1,~\mbox{and}~n\leq\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n-1})+1\leq\mathrm{TC}(M)\leq 2n+1.$ (4)
@@
Precisely, if $F_{i}$ is the $i$-th facet of $P$ and $x_{i}\in H^{1}(M;\mathbb{Z}_{2})$ corresponds to $F_{i}$, then $\iota^{*}(x_{i})\in H^{1}(\mathbb{R}P^{n-1};\mathbb{Z}_{2})$ corresponds to $F_{i}\cap\Delta^{n-1}$. Note that $F_{i}\cap\Delta^{n-1}$ is empty then $\iota^{*}(x_{i})=0$. Since $\iota^{*}$ is surjective, we have
@@
Thus, the left inequality of (4) follows. The right inequality of (4) follows from [17, Theorem 4].
@@
Let $M$ be an $n$-dimensional real torus manifold over $P$ with $n>1$ and $\Delta^{n-1}$ a facet of $P$. Suppose $F_{1},\ldots,F_{n}$ are the facets of $P$ such that $\Delta^{n-1}\cap F_{i}\neq\emptyset$ and $\det[\lambda(F_{1})\cdots\lambda(F_{n})]=1$. Then $\mathrm{cat}(M)=n+1$, and
@@
In particular, if $n=2^{s}$, then
$2^{s+1}\leq\mathrm{TC}(M)\leq 2^{s+1}+1.$
@@
We can have $M=M^{\prime}\#\mathbb{R}P^{n}$ following the discussion in Subsection 1.11 in [10]. Then $H^{*}(\mathbb{R}P^{n};\mathbb{Z}_{2})$ is a subring of $H^{*}(M;\mathbb{Z}_{2})$. Then $n+1=\mathrm{cl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})\leq\mathrm{cl}_{\mathbb{Z}_{2}}(M)\leq n+1$, and the left inequality of (5) follows from the observation $\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})\leq\mathrm{zcl}_{\mathbb{Z}_{2}}(M^{n}(P,\lambda))$. The right inequality of (5) follows from [17, Theorem 4].